# Training the Untrainable: Introducing Inductive Bias via Representational Alignment

### v1

@@ 6 Conclusion
We demonstrated that guidance eliminates the failure modes of networks previously considered unsuitable or ineffective for specific tasks. Aligning with another network overcomes these shortcomings by transferring inductive biases—either architectural and knowledge-based, or solely architectural when using an untrained guide. This opens the door to many applications.
For some networks, like fully connected ones, we only overcame the initial obstacle. Further research is needed to refine these into effective vision models, as they avoid immediate overfitting. This may be a matter of scale, optimizers, learning rates, or other aspects of model design and training. Our results suggest practical applications by significantly narrowing the performance gap between vanilla stacked RNNs and Transformers, albeit in small-scale experiments. Given that stacked RNNs are equivalent to single-layer RNNs, most directly to delayed RNNs (Turek et al., 2020), this implies that complex modifications to RNNs may be unnecessary for language modeling. In future, one might be able to focus on the suitability for inference of different architectures with fewer hard constraints from architecture.
Guidance also proved to be a tool with which to discover the possibility of new initializations. At the moment, no known method exists to find better initializations for networks. In some cases, as with the FCNs for vision, guidance can be disconnected after a nominal number of steps, but still goes on to regularize the target network. This strongly implies that an initialization regime for that target with the same regularization exists. This is all that guidance could do in that case. We now need tools to go backwards, given networks which are correctly initialized and networks which are not, discover what that initialization is. This is a much better place to be in. A systematic sweep of targets and guides to look for better initializations should be carried out.
@@
Our method can be used to study representational and functional design of neural networks in new ways to reinforce or reanalyze prior theory of neural network optimization. For example, we can understand distances between architectural components based on which target networks are easier to guide with a particular guide network. We also refine this notion to include a narrow channel through which guidance can occur, the representational similarity. This can serve as a kind of probe. Different representational similarity methods enforce alignment between different properties of the activations of networks.
Characterizing architectural priors and untrainable networks can be made more precise using our methodology. For example, guidance clearly distinguishes between at least three cases. One where the guide does not help. Another where the architecture of the guide helps, but not the knowledge. And one where both the architecture and the knowledge of the guide help. These are different phenomena whose root causes are not understood at present but that could elucidate the relationship between priors and architectures.
Looking into the long-term future, it may be that architecture can be turned into a prior that we add to the loss of a generic fully-connected network, or any convenient and capable-enough network. Architecture would lose its prominence and the computational substrate could be disconnected from the priors we need for learning and inference. Architecture could potentially even be turned into a parameter to be optimized, rather than a hyperparameter.
Acknowledgments
@@
This work was supported by the Center for Brains, Minds, and Machines, NSF STC award CCF-1231216, the NSF award 2124052, the MIT CSAIL Machine Learning Applications Initiative, the MIT-IBM Watson AI Lab, the CBMM-Siemens Graduate Fellowship, the DARPA Artificial Social Intelligence for Successful Teams (ASIST) program, the DARPA Mathematics for the DIscovery of ALgorithms and Architectures (DIAL) program, the DARPA Knowledge Management at Scale and Speed (KMASS) program, the DARPA Machine Common Sense (MCS) program, the United States Air Force Research Laboratory and the Department of the Air Force Artificial Intelligence Accelerator under Cooperative Agreement Number FA8750-19-2-1000, the Air Force Office of Scientific Research (AFOSR) under award number FA9550-21-1-0399, and the Office of Naval Research under award number N00014-20-1-2589 and award number N00014-20-1-2643. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Department of the Air Force or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
@@ Appendix A Appendix Overview
We present additional details of guidance, experiments and analysis, as well as additional results. In Appendix B, we present some limitations of guidance as a method. In Appendix C, we provide additional details for our guidance approach, with a full explanation of centered kernel alignment. In Appendix D, we provide additional details on our architectural designs and network training. In Appendix E, we introduce a new experiment where we feed noise to our guide network rather than real data and compute representational alignment, leading to similarly improved results. This further establishes a transfer of a prior rather than knowledge. In Appendix F, we provide test accuracy metrics over training as a complementary of network performance over training outside of cross-entropy loss. In Appendix G, we include experiments with a new loss function for task loss, mean-squared error, to demonstrate that guidance is agnostic to the task loss function. In Appendix H, we add an additional experiment where we apply guidance with a new metric for representational alignment, representational similarity analysis, rather than CKA.
Appendix B Methodology Limitations
Our guide network supervision through representational alignment has one primary limitation due to increased memory usage during training. Due to saving activations across several layers of the two networks, GPU memory usage increases dramatically. Moreoever, our methodology works better as batch size increases since this allows for better approximation of representational similarity, increasing memory usage even more. Furthermore, including more layers for supervision leads to improved results.
In this paper, we introduce simple techniques to handle memory constraints such as gradient accumulation and gradient checkpointing (Lin et al., 2017). In practice, more memory optimization techniques may become necessary to consider larger untrainable networks. Further work could consider using stronger representational alignment strategies to reduce the number of samples necessary to achieve a strong fit.
Appendix C Methods Overview
We give an overview of guidance in algorithm 1 and highlight crucial changes to base neural network training in either red or blue. We use blue to indicate the collection of network activations and red to indicate the layerwise mapping and representational alignment using a distance metric. This gives an overview of our layer mapping between the target and guide network. Crucially, we find that the simplest layer mapping where we evenly distribute guide network layers across target network layers for supervision obtains strong results.
@@ Appendix C Methods Overview
18:         $\text{index}\leftarrow\min(\text{round}(i\times step),l-1)$
@@
21:     end for
[…]
23:end for
Appendix D Architecture and Training Details
D.1 Architectural Design Details
@@
D.1.1 Image Classification
@@
Shallow FCN: We design a network similar to Deep FCN but only containing 3 blocks where each feedforward linear layer contains 8192 units. This network is considered untrainalbe due to a saturation on the training performance.
Deep ConvNet: We design a deep convolutional network with the same architecture as ResNet-50 (convolutional layers followed by batch normalization) but remove the residual connections. This makes the network untrainable due to the vanishing gradient problem as observed in He et al. (2016a), causing saturation of the loss.
@@
ResNet-18/50: A deep convolutional network with 18/50 convolutional blocks and residual connections. We refer to He et al. (2016a).
We supervise the Deep FCN and Shallow FCN with ResNet-18 and supervise the Deep ConvNet with ResNet-50.
D.1.2 Copy-Paste
Target Networks
RNN: We design a 4-layer RNN with a hidden dimension of 768 units, followed by a fully connected layer. In copy-paste, architectural and algorithmic limitations make RNNs an untrainable architecture fortask. Specifically, RNNs must memorize the input sequence which is difficult, particularly with a padding token. RNNs are generally considered to be unapplicable to the copy-paste task.
Guide Networks
D.1.3 Parity
Target Networks
Transformer: Similar to Bhattamishra et al. (2020), we design a 1 layer transformer encoder network with a hidden dimension of 64 units across 4 attention heads. Transformers have lower accuracy on formal language tasks that require reasoning over a sequence in comparison to traditional sequence models (Hahn & Rofin, 2024). Due to the enormous gap in performance and saturation of performance, we categorize the transformer as untrainable.
Guide Networks
RNN: We include a 1 layer vanilla RNN with a hidden dimension of 64 units. Similar to Bhattamishra et al. (2020), we achieve 100% accuracy on the task.
D.1.4 Language Modeling
Target Networks
RNN: We design a 5 layer vanilla RNN with a hidden dimension of 512 and with a tanh activation function. We train this with on sequences with a context length of 75. This makes the network untrainable due to problems associated with exploding and vanishing gradients during backpropagation through time.
Guide Networks:
Transformer: We design a 4 layer transformer decoder network with 16 attention heads and a hidden dimension of 512. We train the transformer on WikiText with a context length of 256 and achieve a final test perplexity of 34.15.
D.2 Training
@@
Untrained ResNet-18 $\rightarrow$ Deep FCN $1\times 10^{-4}$
@@
Copy-Paste RNN $1\times 10^{-4}$
Transformer $1\times 10^{-4}$
Transformer $\rightarrow$ RNN $1\times 10^{-5}$
Untrained Transformer $\rightarrow$ RNN $1\times 10^{-5}$
Parity Transformer $1\times 10^{-3}$
RNN $1\times 10^{-2}$
RNN $\rightarrow$ Transformer $1\times 10^{-3}$
Untrained RNN $\rightarrow$ Transformer $1\times 10^{-3}$
Language Modeling RNN $1\times 10^{-4}$
Transformer $1\times 10^{-4}$
Transformer $\rightarrow$ RNN $5\times 10^{-5}$
Untrained Transformer $\rightarrow$ RNN $5\times 10^{-5}$
@@
For all image classification tasks, we used the Adam optimizer (Kingma, 2014), in-line with prior work (He et al., 2016a). For all sequence modeling tasks, we use AdamW (Loshchilov, 2017), which has been useful in training sequence models like RNNs and Transformers (Radford et al., 2019).
@@
Appendix E Representational Regularization
@@
We also aim to understand the role of the guide network as in guidance. In all experiments, we use trained and untrained guide networks and see consistent improvements for training the target network. The success of untrained networks implies that our training method is not performing distillation but instead truly transferring a prior from the guide network to the target network. To more strictly test this theory, we include an experiment where we feed noise to the guide network instead of the same batch of data fed to the target network as implied by eq. 1.
We apply this experiment to the Deep FCN with an untrained ResNet-18 as the target network. At each training step, we pass a noisy batch which is sampled from a random Gaussian with mean of 0 and standard deviation of 1. We train for 100 epochs and report the learning curve results in fig. 6.
@@
Appendix F Test Accuracy across Training

### v2

@@ 6 Conclusion
We demonstrated that guidance eliminates the failure modes of networks previously considered unsuitable or ineffective for specific tasks. Aligning with another network overcomes these shortcomings by transferring inductive biases—either architectural and knowledge-based, or solely architectural when using an untrained guide. This allows guidance to distinguish tasks and architectures that are dependent on architectural biases rather than learned biases. We provide further explanations and intuition for guidance in appendix˜I. There are many potential aspects that may distinguish between the effects of architectural and inductive biases in improving performance of these architectures, which we aim to explore in future work.
@@
This also opens the door to many applications. Our method can be used to study representational and functional design of neural networks in new ways to reanalyze prior theory of neural network optimization. For example, we can understand distances between architectural components based on which target networks are easier to guide with a particular guide network. We also refine this notion to include a narrow channel through which guidance can occur, the representational similarity. This can serve as a kind of probe.
We believe guidance also reveals a new conceptual lens on inductive biases and architectures. When we use an untrained guide network, we reveal what an architecture by itself brings to the table. Guidance with a trained guide reveals how much learned representations can change performance. These are different phenomena whose root causes are not understood at present but that could elucidate the relationship between priors and architectures.
Our results suggest practical applications by significantly narrowing the performance gap between vanilla stacked RNNs and Transformers and showing signs of scaling, albeit in small-scale experiments with 150M parameters or fewer. Given that stacked RNNs are equivalent to single-layer RNNs, most directly to delayed RNNs [73], this implies that complex modifications to RNNs may be unnecessary for language modeling. For other networks, like fully connected ones, we only overcame the initial obstacle. Further research is needed to refine these into effective vision models, as they avoid immediate overfitting. In the future, we hope to find methods for making these networks competitive with their guide networks.
Guidance also proved to be a tool with which to discover the possibility of new initializations. At the moment, no known method exists to find better initializations for networks. In some cases as with the FCNs for vision, guidance can be disconnected after a nominal number of steps, but still goes on to regularize the target network. This strongly implies that an initialization regime for that target with the same regularization exists. This is all that guidance could do in that case. We now need tools to go backwards, given networks which are correctly initialized and networks which are not, discover what that initialization is. This is a much better place to be in. A systematic sweep of targets and guides to look for better initializations should be carried out.
Looking into the long-term future, guidance invites us to treat architecture itself as a trainable prior added directly to a generic network’s loss. Because guidance can rescue models that previously overfit or underfit, we can revisit designs abandoned during neural architecture search. These threads demonstrate the major possibility for future work in this space. Guidance is a tool, not a finished and well-understood theory or doctrine. Tools are useful for enabling more discoveries. We believe that these discoveries will allow the community to more easily pursue questions that training difficulties might have prevented in the past, especially with regard to understanding the relationships between architectures.
Acknowledgements
@@
This work was supported by the Center for Brains, Minds, and Machines, NSF STC award CCF-1231216, the Brains, Minds, and Machines Summer School, the NSF award 2124052, the MIT CSAIL Machine Learning Applications Initiative, the MIT-IBM Watson AI Lab, the CBMM-Siemens Graduate Fellowship, the DARPA Mathematics for the DIscovery of ALgorithms and Architectures (DIAL) program, the DARPA Knowledge Management at Scale and Speed (KMASS) program, the DARPA Machine Common Sense (MCS) program, the Department of the Air Force Artificial Intelligence Accelerator under Cooperative Agreement Number FA8750-19-2-1000, and the Air Force Office of Scientific Research (AFOSR) under award number FA9550-21-1-0399. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Department of the Air Force or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein. V.S. and D.M. are supported by the National Science Foundation Graduate Research Fellowship under Grant No. 2141064. Any opinion, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation.
@@ Appendix A Appendix Overview
We present additional details of guidance, experiments and analysis, as well as additional results. In Appendix˜B, we provide additional details for our guidance approach, with a full explanation of centered kernel alignment. In Appendix˜C, we provide additional details on our architectural designs and network training. In Appendix˜D, we introduce a new experiment where we feed noise to our guide network rather than real data and compute representational alignment, leading to similarly improved results. This further establishes a transfer of a prior rather than knowledge. In Appendix˜E, we show visualizations of the representational dissimilarity loss over training to give context of dynamics over training and show additional explanations for results with trained and randomly initialized guides. In Appendix˜F, we provide further explanation of error consistency as a measure of functional similarity between networks. In Appendix˜G, we provide test accuracy metrics over training as a complementary of network performance over training outside of cross-entropy loss. In Appendix˜H, we provide a comparative baseline to guidance, distillation [33]. We find that basic distillation performs worse in comparison. In Appendix˜I, we provide an interpretation and analysis of guidance to better characterize and understand results based on changes in the internal geometry of a network before and after guidance. In Appendix˜J, we apply guidance with additional neural distance functions including RSA [46] and ridge regression. We find a correlation between the success of our approach and degrees of freedom of a particular distance function. In Appendix˜K, we provide additional ablations over guidance such as guiding a certain number of layers or guiding specific layers of a network.
Appendix B Methods Overview
We give an overview of guidance in algorithm˜1 and highlight crucial changes to base neural network training in either red or blue. We use blue to indicate the collection of network activations and red to indicate the layerwise mapping and representational alignment using a distance metric. This gives an overview of our layer mapping between the target and guide network. Crucially, we find that the simplest layer mapping where we evenly distribute guide network layers across target network layers for supervision obtains strong results.
@@ Appendix B Methods Overview
18:    $\text{index}\leftarrow\min(\text{round}(i\times step),t-1)$
@@
20:    $\text{total}\leftarrow\text{total}+\text{rep}$
21:  end for
22:  $\text{loss}\leftarrow\text{loss}+\text{total}$
23:end for
B.1 Centered Kernel Alignment
To compare representations, we use a representation similarity metric, $\mathcal{M}$, which corresponds to centered kernel alignment (CKA) [45, 16, 17] in our setting. We specifically consider linear CKA.
${\bm{K}}={\bm{R}}{\bm{R}}^{T},{\bm{L}}={\bm{R}}^{\prime}{\bm{R}}^{\prime T}$ (2)
$\tilde{{\bm{K}}}={\bm{H}}{\bm{K}}{\bm{H}},\tilde{{\bm{L}}}={\bm{H}}{\bm{L}}{\bm{H}}$ (3)
We compute the HSIC on the Gram matrices.
@@ B.1 Centered Kernel Alignment
Finally, we define our linear CKA metric as:
$\bar{\mathcal{M}}({\bm{R}},{\bm{R}}^{\prime})=1-\mathcal{M}({\bm{R}},{\bm{R}}^{\prime})$ (6)
Linear CKA ranges from 0 (identical representations) to $1$ (very different representations). Because of this, we take the complement by subtracting the linear CKA from $1$ to represent dissimilarity.
B.2 Methodology Limitations
Our guide network supervision through representational alignment has one primary limitation due to increased memory usage during training. Due to saving activations across several layers of the two networks, GPU memory usage increases dramatically. Moreover, our methodology works better as batch size increases since this allows for better approximation of representational similarity, increasing memory usage even more. Furthermore, including more layers for supervision leads to improved results.
In this paper, we introduce simple techniques to handle memory constraints such as gradient accumulation and gradient checkpointing [51]. In practice, more memory optimization techniques may become necessary to consider larger untrainable networks. Further work could consider using stronger representational alignment strategies to reduce the number of samples necessary to achieve a strong fit.
Appendix C Architecture and Training Details
C.1 Architectural Design Details
@@
C.1.1 Copy-Paste
Target Networks
RNN: We design a 4-layer RNN with a hidden dimension of 768 units, followed by a fully connected layer. In copy-paste, architectural and algorithmic limitations make RNNs an untrainable architecture for the task. Specifically, RNNs must memorize the input sequence which is difficult, particularly with a padding token. RNNs are generally considered to be inapplicable to the copy-paste task.
Guide Networks
C.1.2 Parity
Target Networks
Transformer: Similar to Bhattamishra et al. [8], we design a 1 layer transformer encoder network with a hidden dimension of 64 units across 4 attention heads. Transformers have lower accuracy on formal language tasks that require reasoning over a sequence in comparison to traditional sequence models [28]. Due to the enormous gap in performance and saturation of performance, we categorize the transformer as untrainable.
Guide Networks
RNN: We include a 1 layer vanilla RNN with a hidden dimension of 64 units. Similar to Bhattamishra et al. [8], we achieve 100% accuracy on the task.
C.1.3 Language Modeling
We include two language model settings to test scaling in RNNs. The first setting uses small networks, which we refer to as Small RNN and small Transformer. The other uses a large RNN and large Transformer.
Target Networks
Small RNN: We design a 4 layer vanilla RNN with a hidden dimension of 512 and with a ReLU activation function. We train this on sequences with a context length of 75. This makes the network untrainable due to problems associated with exploding and vanishing gradients during backpropagation through time.
Large RNN: We design a 6 layer vanilla RNN with a hidden dimension of 1024 and with a ReLU activation function. We train this on sequences with a context length of 128. Prior work has demonstrated that larger RNNs are difficult to train in practice [49].
Guide Networks:
Small Transformer: We design a 4 layer transformer decoder network with 16 attention heads and a hidden dimension of 512. We train the transformer on WikiText-103 with a context length of 256 and achieve a final test perplexity of 34.15.
Large Transformer: We design a 4 layer transformer decoder network with 16 attention heads and a hidden dimension of 1024. We train the transformer on WikiText with a context length of 256 and achieve a final test perplexity of 33.10.
C.1.4 Image Classification
@@
Wide FCN: We design a network similar to Deep FCN but only containing 3 blocks where each feedforward linear layer contains 8192 units. This network is considered untrainable due to a saturation on the training performance.
Deep ConvNet: We design a deep convolutional network with the same architecture as ResNet-50 (convolutional layers followed by batch normalization) but remove the residual connections. This makes the network untrainable due to the vanishing gradient problem as observed in He et al. [31], causing saturation of the loss.
@@
ResNet-18/50: A deep convolutional network with 18/50 convolutional blocks and residual connections. We refer to He et al. [31].
We supervise the Deep FCN and Wide FCN with ResNet-18 and supervise the Deep ConvNet with ResNet-50.
C.2 Training
@@
Copy-Paste RNN $1\times 10^{-4}$
Transformer $1\times 10^{-4}$
Transformer $\rightarrow$ RNN $1\times 10^{-4}$
Untrained Transformer $\rightarrow$ RNN $1\times 10^{-4}$
Parity Transformer $1\times 10^{-3}$
RNN $1\times 10^{-2}$
RNN $\rightarrow$ Transformer $1\times 10^{-3}$
Untrained RNN $\rightarrow$ Transformer $1\times 10^{-3}$
Language Modeling Small RNN $1\times 10^{-4}$
Small Transformer $1\times 10^{-4}$
Small Transformer $\rightarrow$ Small RNN $1\times 10^{-4}$
Untrained Small Transformer $\rightarrow$ Small RNN $1\times 10^{-4}$
Large RNN $1\times 10^{-4}$
Large Transformer $1\times 10^{-4}$
Large Transformer $\rightarrow$ Large RNN $1\times 10^{-4}$
Untrained Large Transformer $\rightarrow$ Large RNN $1\times 10^{-4}$
@@
Untrained ResNet-18 $\rightarrow$ Deep FCN $5\times 10^{-5}$
@@
For all image classification tasks, we used the Adam optimizer [43], in-line with prior work [31]. For all sequence modeling tasks, we use AdamW [52], which has been useful in training sequence models like RNNs and Transformers [61].
@@
Appendix D Representational Regularization: Guidance with Noise
@@
We also aim to understand the role of the guide network as in guidance. In all experiments, we use trained and untrained guide networks and see consistent improvements for training the target network. The success of untrained networks implies that our training method is not performing distillation but instead truly transferring a prior from the guide network to the target network. To more strictly test this theory, we include an experiment where we feed noise to the guide network instead of the same batch of data fed to the target network as implied by eq.˜1.
We apply this experiment to the Deep FCN with an untrained ResNet-18 as the target network. At each training step, we pass a noisy batch which is sampled from a random Gaussian with mean of 0 and standard deviation of 1. We train for 100 epochs and report the learning curve results in fig.˜6.
@@
Appendix E Representational Similarity Loss
Figure 7: CKA representational similarity loss for sequence modeling tasks. We visualize the total CKA dissimilarity loss across all layers across training for all three sequence modeling tasks. The CKA dissimilarity loss represents the representational alignment between our guide network and target network. We can observe that for the copy-paste task and language modeling task, the target network aligns with a randomly initialized network more quickly. This could be because of special properties of RNNs.
Figure 8: CKA representational similarity loss for image classification. We visualize the total CKA dissimilarity loss across all layers across training for image classification. The CKA dissimilarity loss represents the representational alignment between our guide network and target network. We can observe that for Deep FCN and Wide FCN, the target network aligns with a randomly initialized network more quickly. This corresponds with results where randomly initialized guide networks had superior performance to trained guide networks.
We can view the representational alignment between the guide and target networks during training. This allows us to better understand how this representational alignment influences network performance. We show sequence modeling results in fig.˜7 and image classification results in fig.˜8.
We notice that across most tasks, reducing representational dissimilarity is easier with activations from randomly initialized networks rather than trained networks. This provides additional evidence of representational alignment for inductive bias transfer. We notice that for certain cases, such as Parity, the randomly initialized guide network has higher representational dissimilarity loss than the trained guide network. This is matched with the Parity result in table˜2 and fig.˜2.
However, we can also observe more inconsistent results with the Deep ConvNet where the untrained guide network has lower representational dissimilarity loss than the trained guide network, even at the end of training. One possible explanation that the inductive bias was more similar for Deep ConvNet and ResNet-50. This means that trained features are more important for better Deep ConvNet results and representational alignment with a trained network is important. This result has interesting implications for understanding the role of residual connections. Since untrained ResNet-50 is easier to align with than a trained ResNet-50, this demonstrates that residual connections influence representation spaces during training. The untrained residual connections have little influence on the inductive biases of the network or the overall representation space. This demonstrates the strength of our method as a way to interpret neural network design choices and how they influence representation and functional aspects of a network.
These results are also potentially indicative of architectural properties of RNNs and FCNs which match randomly initialized networks more quickly. For instance, one potential explanation is that RNNs have more degrees of freedom [8] and therefore, only need inductive guidance rather than trained features. Transformers may require learned features indicating that the bottleneck for transformers on the parity is not algorithmic but feature-based. Much of the future work can use these results to design better networks with more informed designs.
E.1 Layerwise Analysis
Figure 9: CKA dissimilarity decreases more in earlier network layers than later layers. When we separate the CKA dissimilarity across layers of the target network, we find that earlier layers optimize more and earlier. We take two layers from a Deep FCN during guidance with a randomly initialized ResNet-18. The early layer comes from the 15th FCN block. The later layer comes from the 43rd FCN block. We see that both layers are eventually optimized but the later layer receives less supervision and has a higher CKA at the end of training.
We provide a deeper analysis of patterns of the representational dissimilarity across different layers during guidance in fig.˜9. We find that earlier layers generally have higher CKA similarities with their corresponding layer from the guide network and later layers have lower CKA similarities. Furthermore, these later layers optimize later in the training process.
Appendix F Error Consistency
We measure error consistency $(\kappa)$ [23] between the guided target networks which indicates the error overlap between two networks based on the accuracy of the networks, i.e. do the two networks make similar class predictions? The measure first calculates the expected error overlap. Suppose $a_{1}$ is the accuracy of the first guided network and $a_{2}$ is the accuracy of the second. The expected error overlap is given by $c_{\text{exp}}=a_{1}*a_{2}+(1-a_{1})*(1-a_{2})$. Next, we measure the observed error overlap across each sample in the validation set as $c_{\text{obs}}=\text{\# of samples where both models agree}~/~\text{total trials}$. Finally, we can write $\kappa$ as:
$\kappa=\frac{c_{\text{obs}}-c_{\text{exp}}}{1-c_{\text{exp}}}$ (7)
$\kappa$ ranges from 1-1 to $1$, where $1$ is perfect agreement, 1-1 is perfect disagreement and 0 is change agreement. When $\kappa>0$, this implies that models make consistent error patterns, $\kappa<0$ implies that models make inverse error patterns, and $\kappa\approx 0$ implies independent error patterns.
F.1 Guide Network Representation Comparison
Figure 10: Earlier layers of ResNet-18 and ViT-B are more similar. We analyze the representational similarity between activations from layers ResNet-18 and ViT-B-16 via CKA. We find that earlier layers are more similar while later layers have divergent representations. We see that this manifests in distinct error consistency patterns when ResNet-18 and ViT-B are used as guide networks.
We contextualize the findings in error consistency by comparing the representations of the guide networks, in this case ResNet-18 and ViT-B-16. We apply a layer mapping between layers of ResNet-18 and ViT-B-16 and compute the representational similarity over 1000 input images. Results are shown in fig.˜10.
Our findings are useful for error consistency. If models are inheriting inductive biases from their guide network, then the models would have similar methods to process low-level image features are indicated by a stronger CKA in earlier layers between the ResNet-18 and ViT-B. This means that errors will be consistent for low level features but inconsistent for high level features collected in later layers.
Appendix G Test Accuracy across Training